OpenStudy (chillout):

Show that \((A-I)^{-1}=I+A+A^{2}+A^{3}\) if \(A^{4}=0\)

OpenStudy (vishweshshrimali5):

Try pre-multiplying with (A-I) on both sides...

OpenStudy (chillout):

One second thought, this was not the question! I've done this one. Sorry!

OpenStudy (chillout):

On second thought*

OpenStudy (chillout):

I'm editing it!

OpenStudy (vishweshshrimali5):

Yeah sure!

OpenStudy (vishweshshrimali5):

So what's the new question?

OpenStudy (chillout):

I'm taking a picture. It's a big matrix and I'm supposed to use \(sgn(\sigma)\) to calculate it.

OpenStudy (kainui):

Are you familiar with how to derive the geometric series: \[\frac{1}{1-x} = 1+x+x^2+x^3+x^4+\cdots\]

OpenStudy (vishweshshrimali5):

Kai does that work for matrices as well??

OpenStudy (chillout):

Yeah, I am.

OpenStudy (kainui):

@vishweshshrimali5 It does if you're careful :O Follow the steps you'd take to derive it for scalars but do it in a matrix sorta way. :D

OpenStudy (vishweshshrimali5):

:p Ohh well... basically like starting from beginning :P

OpenStudy (chillout):

I am supposed to use permutations and the signal function, as the Leibniz/Laplace formula

OpenStudy (chillout):

But a tree approach wouldn't work here? Like, all possible multiplications because it's a matrix with a lot of 0's.

OpenStudy (chillout):

Nevermind the Laplace method. Just the Leibniz one.

OpenStudy (vishweshshrimali5):

You have to show the equation for this matrix?

OpenStudy (chillout):

I'm forgetting something about the \(sgn(\sigma)\) function.

Parth (parthkohli):

What are these words? Don't we just have to evaluate the determinant?

OpenStudy (vishweshshrimali5):

Just use alternative positive and negative coeff. that's what signum function is all about

OpenStudy (chillout):

I'm supposed to evalute using the Leibniz method! But might as well might find an easier way.

OpenStudy (vishweshshrimali5):

Wait! Let's go for Leibniz if it's specified

OpenStudy (chillout):

@vishweshshrimali5, the wiki article isn't much of help.

Parth (parthkohli):

Convert it to a diagonal matrix. You'll just need two row operations. Not familiar with the other one though.

OpenStudy (vishweshshrimali5):

\[\det(A)=\sum_{\sigma \in S_n}\text{sgn}(\sigma)\prod_{i=1}^{n} a_{\sigma(i),i}\]

OpenStudy (vishweshshrimali5):

Here we have 5 by 5 matrix

OpenStudy (vishweshshrimali5):

So there will be 5! combinations

OpenStudy (chillout):

Am I wrong or there will be a lot of operations?

OpenStudy (vishweshshrimali5):

For signum: We have to see the oddness/evenness of the sequence...

OpenStudy (vishweshshrimali5):

There will be a big lot but because we have so many zeroes... many should become zero

OpenStudy (vishweshshrimali5):

Now, evenness or oddness can be defined as follows: the permutation is even (or odd) if the new sequence can be obtained by an even number (odd, respectively) of switches of numbers. Right?

OpenStudy (chillout):

Hm... Sounds okay, but how about the number of switches? I'm assuming this is not arbitraty otherwise this method would fail.

OpenStudy (vishweshshrimali5):

Naah it's not arbitrary ... we find out the evenness/oddness w.r.t [1,2,3,4,5] set

OpenStudy (vishweshshrimali5):

Anyways so if I write the det using Leibniz formula: \[det(A) = (-1)^0 a_{1,1} a_{2,2} a_{3,3} a_{4,4} a_{5,5} + (-1)^1 a_{1,1} a_{2,2} a_{3,3} a_{4,5} a_{5,4} + ...\]

OpenStudy (vishweshshrimali5):

Now only non zero terms are: \(a_{1,1}, a_{1,4}, a_{2,2}, a_{2,5}, a_{3,3}, a_{4,4}, a_{5,4}, a_{5,5}\)

OpenStudy (vishweshshrimali5):

So any combination having terms outside of the above will become zero....

OpenStudy (vishweshshrimali5):

This definitely makes it easier... doesn't it?

OpenStudy (chillout):

Affirmative. So \(sgn(\sigma)\) just grow by 1 for every switch?

OpenStudy (vishweshshrimali5):

The power of (-1) goes up by 1 for every switch

OpenStudy (vishweshshrimali5):

sgn will have only two possible values -1 or 1

OpenStudy (vishweshshrimali5):

But remember there will be 5! combinations = 120 ... so instead of trying to write the whole series... just use those 8 non zero terms to get the possible non zero combinations

OpenStudy (chillout):

I'm getting confused in the permutation. I'm not sure how to find if it's even or odd. For example, the set \({a,b,c,d,e}\) switched to \(a,c,b,e,d\) has which kind of parity?

OpenStudy (vishweshshrimali5):

Actually this is where I am getting confused too :P Never used Leibniz formula so have no idea... will have to look into it... @Kainui any ideas?

Parth (parthkohli):

Just do \(R_5 \to R_5 - \frac{3}5 R_4\) followed by \(R_2 \to R_2 - \frac{1}2 R_5\). Now you're left with a diagonal matrix. And it's an obvious fact that for a diagonal matrix, the determinant is just the product of diagonal entries, which are \(2, 3, 4, 5, 2\).

OpenStudy (kainui):

I think they're expecting you to do this fast and use the Laplace expansion to expand along every row with zeroes in it except for one of the entries. The sign of the permutation doesn't matter for a 0 entry.

OpenStudy (vishweshshrimali5):

Yeah ... Leibniz is looking too messy to be used here...

OpenStudy (kainui):

Well, you can definitely use Leibniz, but don't spend time counting the 100 some odd terms that are all zero cause no matrix is worth that much time of your life haha

OpenStudy (chillout):

Actually the question asks to use the definition of determinant, but Laplace is shown shortly after it. That's why I took it back when I said I could use Laplace :P

OpenStudy (vishweshshrimali5):

And there comes the crux of the problem by our very own Kainui ;) Even if we go for Leibniz, we can directly use the 8 non negative terms to make our work easier... but it will be the sgn that will be a pain to find out..

OpenStudy (vishweshshrimali5):

If you have read about elementary row operations then as @ParthKohli showed just use them... they are much easier here....

OpenStudy (chillout):

I've figured as much. They wouldn't give a matrix with lots of numbers \(\neq\)0

OpenStudy (vishweshshrimali5):

Yeah...

OpenStudy (chillout):

But the definition of determinant is so vague I.M.O.

OpenStudy (kainui):

I already had this typed up and didn't want to interrupt, but I am getting antsy so I'm going to drop this now. Ignore it and come back to it later please cause I want to talk about the determinant stuff now. ---- First let's start with this infinite sum. We can do the finite case and get the same answer, but since \(A^4=0\) it really isn't worth it imo \[S=I+A+A^2+\cdots\] Now we multiply both sides by \(A\) \[AS=A+A^2+A^3+\cdots\] Add \(I\) to both sides: \[AS+I = I+A+A^2 +\cdots = S\] Now we add and subtract to get S on one side \[S-AS=I\] Factor S out \[(I-A)S=I\] We were specifically told to prove something about \((I-A)^{-1}\) so it definitely exists, thus we can left multiply both sides of this equation by it to get what we always wanted deep in our hearts: \[S=(I-A)^{-1}\] Since S is that and it's also that sum, which is what we wanted to prove, \[(I-A)^{-1}=I+A+A^2+A^3\] We're done! If you don't believe this infinite thing you can walk through the exact same stuff with the finite sum and when you see \(A^4\) kill it. Also if you want you can add \(A^4\) when you need it if necessary, cause it's 0 and you can add 0 to anything and not hurt it.

OpenStudy (vishweshshrimali5):

Well I can't say that :P My whole work revolves around dealing with matrices, determinants for data storage and computation :p

OpenStudy (vishweshshrimali5):

Okay @Kainui _/\_ Wow!!

OpenStudy (chillout):

woah. Actually there's some problem in the set for \(A^{n+1}\). You've answered it too :P

OpenStudy (kainui):

so what definitions of determinants are you given and expected to work this problem with haha

OpenStudy (vishweshshrimali5):

Well guys I gtg... have an early class to attend tomorrow :(

OpenStudy (vishweshshrimali5):

Till then.. keep solving :P Good night :)

OpenStudy (kainui):

Nice seeing you around vishwesh, it's been a while :P good night

OpenStudy (chillout):

I was given Leibniz formula. Laplace method is given shortly afterwards :(

OpenStudy (chillout):

Thanks @vishweshshrimali5 for your effort!

OpenStudy (kainui):

Alright so does your book use this exact notation or something else, I want to be consistent with your book and class instead of some wikipedia article: \[\det(A)=\sum_{\sigma \in S_n}\text{sgn}(\sigma)\prod_{i=1}^{n} a_{\sigma(i),i}\]

OpenStudy (chillout):

Yeah that's the notation.

OpenStudy (chillout):

I'm having trouble with the permutations

OpenStudy (chillout):

For example, let \(S=\{1,2,3\}\). It says that (3,2,1) has 3 switches!

OpenStudy (kainui):

So \(\sigma\) represents the numbers 1 through 5 in a specific order, like (3,5,2,1,4). Since \(\det(A)=\det(A^\top)\) we don't have to worry about talking about rows or columns of a determinant, I'll just talk about rows to explain what that example above means. It means we pick the 3rd entry from the first row, then the 5th entry from the second row, then the 2nd entry on the 3rd row, the 1st entry off of the 4th row and the 4th entry from the last row. These are what that little indexed sigma means: \(\sigma(i)\) so in my example, \(\sigma(2)=5\). Kinda a lot going on, but the main takeaway here is that this severely limits the nonzero products you can pick from the matrix, so we have two choices for the first row we can pick from the first or 4th entry on that row. I'll go ahead and pick the 4th one in the row, it's a 2. Next I look at the second row, I have a choice between a 3 and 1, I'll pick the 3 why not. Then I get to the 3rd row and I have only one option. So I pick the 4 and keep going to the next row where I have a 5. It's in the 4th entry of the 4th row and it's the ONLY entry. But my very first decision was something from the 4th entry of the 1st row, so I'm forced to pick a 0 entry if I start with it. So here's what my permutation would have looked like: \((4,2,3, 4, \ )\) I don'tknow if that makes it better or worse

OpenStudy (kainui):

For example, let \(S=\{1,2,3\}\). It says that (3,2,1) has 3 switches! Yes, so every permutation has a "standard order" and this one's is (1,2,3). Apparently they're limiting you to only doing switches between adjacent pairs, although technically all that really matters is the parity of the switch which is the same if you just immediately flip the 1 and 3, that's an odd number of flips. The way they're doing it is one of two possible ways: (3,2,1) -> (3,1,2) -> (1,3,2) -> (1,2,3) idk why they're making you do flips between adjacent ones only, but it's more concrete I suppose since this is all pretty whacked out and they're trying not to confuse you... So if I've confused you, well, happens I guess lol.

OpenStudy (chillout):

This method is so confusing. I might leave it aside because... Just because. I'm grateful for your answers and your help but I am unlikely to revisit this method ever again, lol. Back to elementary rows and Laplace... :(

OpenStudy (kainui):

Yeah, this method is terrible I agree haha. The only upside of the Leibniz formula is that it's a formula that you can use in equations... Which is still pretty disgusting, but at least it can be used in higher level physics to "plug and chug" determinants with stuff if needed ahaha